OpenStudy (anonymous):

the x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line?

OpenStudy (anonymous):

Do you know the standard line equation?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

There are two ways to approach the problem we can use y=mx+b and solve for m. Or we can use m=(y2-y1)/(x2-x1). Which are you learning?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

well perhaps the easiest way is to just use the slope equation. The x intercept is when y=0, thus point one (x1,y1)=(-5,0). The y intercept is when x =0. Thus point two (x2,y2)=(0,-2). Now use the equation I wrote above to solve for m. Do you knwo what to do once you find m?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well, as I mentioned above, the standard equation of a line is in the form of y=mx+b. YOu are now solving for m, and the y intercept is b. thus b=2

OpenStudy (anonymous):

b=-2

OpenStudy (anonymous):

so i use m=y2-y1/ x2-x1?

OpenStudy (anonymous):

yes, make sure you use your order of operations. (y2-y1)/(x2-x1)

OpenStudy (anonymous):

I got -2/5?

OpenStudy (anonymous):

(-2-0)/(0-(-5)) =-2/5.. yes you are correct